A thin rod 9m 6cm long is
cut into two so that one piece is 96cm longer than the other. What is the length of the longer piece

Answer by LinnW(1048)

The original rod is 9m 6cm long = 900 + 6 cm long = 906 cm
Let x = the shorter length
x + 96 = the longer length
The shorter length plus the longer length = 906
x + (x + 96) = 906
2x + 96 = 906
add -96 to each side
2x = 906 - 96
2x = 810
x = 405 which is the shorter length
x + 96 = 405 + 96 = 501 which is the longer length